# Is deeper always better? Replacing linear mappings with deep learning networks in the Discriminative Lexicon Model

### v1

@@ 4.1 Predicting average reaction times
For both the DLP and BLP we trained 5 comprehension models:
@@
We used CELEX (Baayen et al.,, 1995) frequencies for both the BLP and DLP throughout. After training, we obtained the correlation between the predicted and the target semantic vectors for each model (“target correlation”). Reaction times were transformed as follows:
@@
We used the target correlation measure to predict RTinv in one GAM per model. Target correlation was a highly significant predictor for reaction times across all models and both datasets ($p<0.0001$; summary tables for the GAMs can be found in the Supplementary Materials). Figure 4 shows the fit of the 5 models in terms of the Akaike Information Criterion (AIC, Akaike,, 1998) for both the DLP and BLP. First, we see that for both datasets, frequency-informed methods (FIL, FIDDL) outperform their non-frequency-informed counterparts (EL, DDL, EDDL). Regarding the difference between DDL and LDL, DDL and EDDL outperform EL for the DLP, but not for the BLP. Interestingly, while the correlation accuracy for EDDL was clearly higher than for DDL (DLP: 99% for EDDL compared to 92% for DDL; BLP: 100% compared to 88%), when it comes to predicting reaction times, EDDL only shows a slight advantage over DDL for the BLP and performs worse for the DLP. This indicates that better mapping accuracy does not necessarily entail better performance at predicting behavioural data. FIDDL outperforms all other models. From this we can conclude that while DDL does not necessarily outperform LDL, FIDDL clearly outperforms LDL both without and with frequency information. We will return to both these points in the final discussion.
(a) BLP
(b) DLP
Figure 4: Target correlation taken from an EL, FIL, DDL, EDDL and FIDDL model predicting reaction times in the BLP and DLP. Figure adapted from Heitmeier, (2024).
@@ 4.2 Predicting trial-to-trial reaction times
Next, we turn to using DDL models for predicting more fine-grained data, namely lexical decision data at the per-participant trial-to-trial level. Heitmeier et al., 2023b hypothesised that humans’ lexical knowledge is continuously changing and that this continuous change can be modelled with the DLM. They simulated the lexical decision experiment reported in the BLP. For each participant they ran one simulation where they updated the LDL mappings in the DLM using the Widrow-Hoff (Widrow and Hoff,, 1960) learning rule (dynamic simulation) and one where they did not update the mappings (static simulation). After applying one learning step of the Widrow-Hoff learning rule, the error between a predicted and target vector is slightly lower than before the update. This allows updating the mappings after each encountered stimulus in the lexical decision experiment in our simulations (details in Heitmeier et al., 2023b, ; Heitmeier,, 2024). Measures from these simulations were then used to predict individual reaction times in the BLP. They found that for a majority of subjects, dynamic simulations provided a better model fit to reaction times, particularly for words (as opposed to nonwords).
@@
Here, we asked whether we can replicate these findings with DDL. To this end we repeated the analysis presented in Heitmeier et al., 2023b for the first 39 subjects in the BLP, replacing LDL mappings with DDL models and applying one backpropagation step (instead of the Widrow-Hoff learning rule) for updating the models after each trial. Details on the simulation procedure, the measures used for predicting reaction times and the statistical modelling can be found in Heitmeier, (2024).
@@
We found that the AIC of all DDL-based models was worse than that of the LDL-based counterparts for all participants, for both words and nonwords and static/dynamic simulations (average AIC difference was -109.0/-131.6 AIC points for words for static/dynamic simulations and -126.1/-119.6 AIC points for nonwords for static/dynamic simulations). While for LDL we found that 72% of dynamic simulations showed a better fit than static simulations for words (33% for nonwords), for DDL 67% of dynamic simulations outperformed static ones for words (46% for nonwords). Figure 5 gives an overview over the average difference in AIC as well as the percentage of participants showing an advantage of trial-to-trial learning.
@@
Figure 5: Comparison between static and dynamic simulations using DDL and LDL for predicting per-participant reaction times in the BLP. In general, the effect of dynamic simulations compared to static ones is larger for LDL, resulting in a bigger difference in AIC as well as a larger number of participants showing an advantage of trial-to-trial learning for LDL. Figure from Heitmeier, (2024).
@@
When plotting the difference between the AICs of static and dynamic simulations with DDL against that with LDL (Figure 6), we found that they correlate for words, i.e. DDL and LDL generally agreed on which subjects show an effect of trial-to-trial learning as modelled by the dynamic simulation. No such correlation was present for nonwords.
@@
Figure 6: Comparison of AIC differences between dynamic and static simulations for DDL and LDL. Each dot represents one participant. The x-axis denotes the difference in AIC between dynamic and static simulations for the LDL models and the y-axis for the DDL models. Negative difference values indicate that the dynamic simulation outperforms the static simulation in terms of AIC, while positive difference values indicate the opposite. If a dot falls within a green shaded area, LDL and DDL agree on whether dynamic simulations provide a better model fit that static simulations for the respective participant. For words, the AIC differences are correlated ($r=.51,p=0.001$); for words no correlation is present ($p=0.436$). Figure from Heitmeier, (2024).
We conclude from this analysis that DDL is not able to model fine-grained trial-to-trial effects as well as LDL. DDL shows an advantage of learning for fewer subjects than LDL and the difference between dynamic and static is smaller for DDL than for LDL. However, at least for words, DDL and LDL largely agree on which subjects show an effect of trial-to-trial learning. A possible reason for the reduced effect of trial-to-trial learning in DDL is that a more accurate initial mapping may lead to less pronounced error feedback in the update step, thus each learning step having a smaller impact on subsequent trials. Moreover, the larger number of weights in the DDL networks might also mean that weight updates are distributed more evenly across weights and not centered as strongly on the weights of cues shared with subsequent words.
@@ 5 Discussion and Conclusion
In this paper we explored whether replacing the linear mappings (Linear Discriminative Learning, LDL) in the Discriminative Lexicon Model (DLM) with dense neural networks (termed Deep Discriminative Learning, DDL) improves the model’s mapping accuracy as well as its performance at predicting behavioural data. We found that across four languages (English, Dutch, Taiwan Mandarin, Estonian) DDL always improved training accuracy and also accuracy on held-out data for English and Dutch, the two datasets with the largest number of different lemmas. When predicting average lexical decision reaction times, we found that a frequency-informed variant of DDL (FIDDL) where the model is trained on each word according to its frequency outperformed all other methods (including frequency-informed LDL, FIL). However, when predicting fine-grained trial-to-trial lexical decision reaction times, DDL performs worse than LDL, even though the two generally agree as to which subjects show a better model fit when based on simulations where the DLM is dynamically updated after each trial.
@@
It is worthwhile to focus on a few aspects of these results in more detail. First, we find a notable difference across datasets: In comprehension, DDL only reliably improves mapping accuracy over that of LDL for English and Dutch, but not for Mandarin and Estonian. This may to some extent be explainable by the respective data sizes: the English and Dutch datasets are larger (28,000 and 11,000 word forms) than the Estonian and Mandarin ones (6,000 and 3,000). However, the nature and diversity of the datasets may also play role. The English and Dutch datasets have a much larger number of different lemmas with few inflected word forms per lemma, while the Estonian dataset only has 230 lemmas with a large number of (relatively regular) inflected forms. For Mandarin, tone information contributes to a much larger number of cues (9,644, compare this to the much larger English dataset, which only has 7,030 cues), and as we saw in Section 3, fewer shared cues increase the performance of LDL models in comprehension. This suggests that the more semantically diverse a dataset, the more it profits from modelling with DDL. In production, on the other hand, all four models profit from DDL. In Heitmeier, (2024) we found that this can be accounted for to a significant extent by the loss function used: a linear model with the same number of trainable parameters as LDL but with added sigmoid + binary cross entropy resulted in almost the same accuracy as the deep learning models. However, the advantage of DDL over LDL in the present setup may also be due to the representations used. Chuang et al., (2024) found that LDL performs surprisingly well at mapping contextualised embeddings to pitch contours, and preliminary results suggest that in this setup, DDL does not improve over LDL even in production. This indicates that when both input and output are sufficiently discriminable, LDL suffices.
@@
Second, notice that while DDL models outperformed LDL models in terms of pure mapping accuracy for both the English and Dutch datasets, they do not necessarily outperform LDL at predicting reaction times: DDL outperforms LDL at predicting average reaction times in the Dutch Lexicon Project (DLP), but not in the British Lexicon Project (BLP). The same is true for modelling lexical decision reaction times at the individual subject-level: even leaving aside trial-to-trial learning, the DDL models never perform better than the LDL models for static simulations. This highlights that pure mapping accuracy is not enough for predicting reaction times. However, it also leads to a further point, namely that modern machine learning and artificial intelligence may in some aspects be too good: human speakers make a lot of mistakes and are not particularly accurate. For example, participants in the DLP agree with themselves at chance level for low frequency words (below 10 per million) when tested on the same words twice (Diependaele et al.,, 2012). In the study of Ramscar et al., (2013), young adults (mean age 21.1 years) performed at chance level for low frequency words in an English lexical decision task (Balota et al.,, 1999). Moreover, the question arises of whether fully optimized performance in the DLM would be in any way helpful for predicting reaction times. Dense neural networks can in principle learn any continuous function to an arbitrary point of accuracy when given enough parameters (Hornik et al.,, 1989). In the case of the DLM this would mean that all predicted semantic and form vectors would be entirely accurate, thus losing any information we can gain from the DLM above and beyond simply using binary form vectors and pre-trained embeddings. The DLM’s imprecisions and mistakes offer valuable insights into possible processing bottlenecks.
Interestingly, from an evolutionary perspective, there is also some evidence that natural selection may not always favour the most optimal rule: Trimmer et al., (2012) found that in a simple evolutionary simulation, the Rescorla-Wagner learning rule was often favoured over other, more optimal learning rules, since it is more robust to variations in its parameterisation. For our setup this could suggest that even though deep learning models are more accurate, the higher number of parameters may make it less robust compared the simpler linear networks.
@@
Third, examining the words that DDL maps more accurately than LDL is illuminating for linguistic analysis. We found that LDL outperforms DDL for words which contain very unique trigrams (e.g. #px, #tv), while DDL outperforms LDL when words contain trigrams which express a range of different functions such as -er. The re-use of building blocks is ubiquitous in language. While humans are able to deal with this “bricolage” relatively effortlessly, it is nevertheless a processing bottleneck which LDL struggles with. In this way, the comparison between LDL and DDL shows which components of single word processing may be particularly challenging. This finding could also have implications for theories of morphological priming (e.g. Taft and Forster,, 1975; Rastle and Davis,, 2008): our results suggest that LDL may place pseudomorphological words (e.g. slender, choler) in the semantic neighbourhood of morphologically complex words (e.g. slighter, chunkier). This could explain why morphologically transparent and morphologically opaque words are primed in a similar way (e.g. Creemers et al.,, 2020) even without morphological decomposition. Again, however, this result may, at least in part, be due to the representations used. An examination of the final segment of words ending in -er in the Buckeye corpus (Pitt et al.,, 2007) reveals that they have more varied pronunciations in nouns than in adjectives. These results dovetail well with the hypothesis that the nouns, which have much more varied meanings, have more differentiated realizations of the exponent — a distinction that is neutralized in the orthography. This in turn suggests that deep learning is required for -er words in English because the representations are impoverished, which opens the possibility that a linear mapping based on actual articulation might suffice.
@@
Fourth, our results highlight the importance of the data distribution the models are trained on for cognitive modelling. Using a training/validation/test split is not a plausible mechanism for training cognitive models: after all, we usually encounter high frequency (seen) words, with the occasional low frequency (possibly unseen) word interspersed in-between. FIDDL (frequency-informed DDL) therefore has a much more plausible learning setup: it is trained on the full token distribution. By calculating the model’s accuracy across all words in the dataset we get a more realistic picture of how the model would fare with a real-world distribution of tokens (see also Heitmeier et al., 2023a, , for a discussion of token-based accuracy calculations). Surprisingly, this accuracy is very high (97% and 87% for Dutch and English respectively) even though the model saw 40% of the words in the English and 9% of the words in the Dutch dataset only 10 times or fewer. Such a setup also quells concerns about the performance of computational models on held-out data (Pinker and Ullman,, 2002). Humans are able generalise to unseen data, but in reality, word distributions are skewed towards high frequency words. Evidently, both humans and FIDDL perform excellently in such a setting. Moreover, these results once again highlight the importance of training models of lexical processing in a frequency-informed way when one is interested in a usage-based analysis of the data. The frequency-informed FIDDL and FIL models by far outperform their non-frequency informed counterparts when predicting average reaction times in the BLP and DLP.

### v2

@@ 4.1 Predicting average reaction times
For both the DLP and BLP we trained five comprehension models:
@@
We used CELEX (Baayen et al.,, 1995) frequencies for both the BLP and DLP throughout. After training, we obtained the correlations between the predicted and the target semantic vectors for each model (“target correlation”). Reaction times were transformed as follows:
@@
We used the target correlation measure to predict RTinv in one GAM per discriminative learning model. Target correlation was a highly significant predictor for reaction times across all models and both datasets ($p<0.0001$; summary tables for the GAMs can be found in the Supplementary Materials). Figure 4 shows the fit of the five models in terms of the Akaike Information Criterion (AIC, Akaike,, 1998) [^6] AIC measures model fit while punishing model complexity. Lower AIC values indicate better model fit. for both the DLP and BLP. First, we see that for both datasets, frequency-informed methods (FIL, FIDDL) outperform their non-frequency-informed counterparts (EL, DDL, EDDL). Regarding the difference between DDL and LDL, DDL and EDDL outperform EL for the DLP, but not for the BLP. Interestingly, while the correlation accuracy for EDDL was clearly higher than for DDL (DLP: 100% for EDDL compared to 93% for DDL; BLP: 100% compared to 89%), when it comes to predicting reaction times, EDDL only shows a slight advantage over DDL for the BLP and performs worse for the DLP. This indicates that better mapping accuracy does not necessarily entail better performance at predicting behavioural data. FIDDL outperforms all other models. From this we can conclude that while DDL does not necessarily outperform LDL, FIDDL clearly outperforms LDL both without and with frequency information. We will return to both these points in the final discussion.
(a) AIC for predicting RTs in the BLP
(b) AIC for predicting RTs in the DLP
Figure 4: Target correlation taken from an EL, FIL, DDL, EDDL and FIDDL model predicting reaction times in the BLP and DLP. Figure adapted from Heitmeier et al., 2025b .
@@ 4.2 Predicting trial-to-trial reaction times
Next, we turn to using DDL models for predicting more fine-grained data, namely lexical decision data at the per-participant trial-to-trial level. Heitmeier et al., 2023b hypothesised that humans’ lexical knowledge is continuously changing and that this continuous change can be modelled with the DLM. They simulated the lexical decision experiment reported in the BLP. For each participant they ran one simulation where they updated the LDL mappings in the DLM using the Widrow-Hoff (Widrow and Hoff,, 1960) learning rule (dynamic simulation) and one where they did not update the mappings (static simulation). After applying one learning step of the Widrow-Hoff learning rule, the error between a predicted and target vector is slightly lower than before the update. This allows updating the mappings after each encountered stimulus in the lexical decision experiment in our simulations (details in Heitmeier et al., 2023b, ; Heitmeier et al., 2025b, ). Measures from these simulations were then used to predict individual reaction times in the BLP. They found that for a majority of subjects, dynamic simulations provided a better model fit to reaction times, particularly for words (as opposed to nonwords).
@@
Here, we asked whether we can replicate these findings with DDL. To this end we repeated the analysis presented in Heitmeier et al., 2023b for the first 39 subjects in the BLP, replacing LDL mappings with DDL models and applying one backpropagation step (instead of the Widrow-Hoff learning rule) for updating the models after each trial. As in the linear simulations, each backpropagation step included calculating the loss of only this trial’s predicted vector and using this loss to update the model’s weights. Details on the simulation procedure, the measures used for predicting reaction times and the statistical modelling can be found in Heitmeier et al., 2023b and Heitmeier et al., 2025b .
@@
We found that the AIC of all DDL-based models was worse than that of the LDL-based counterparts for all participants, for both words and nonwords and static/dynamic simulations (average AIC difference was -108.5/-128.4 AIC points for words for static/dynamic simulations and -128.3/-118.1 AIC points for nonwords for static/dynamic simulations). While for LDL we found that 72% of dynamic simulations showed a better fit than static simulations for words (33% for nonwords), for DDL 74% of dynamic simulations outperformed static ones for words (62% for nonwords). Figure 5 gives an overview over the average difference in AIC as well as the percentage of participants showing an advantage of trial-to-trial learning.
@@
Figure 5: Comparison between static and dynamic simulations using DDL and LDL for predicting per-participant reaction times in the BLP. For words, the effect of dynamic simulations compared to static ones is larger for LDL, resulting in a bigger difference in AIC. Figure adapted from Heitmeier et al., 2025b .
@@
When plotting the difference between the AICs of static and dynamic simulations with DDL against that with LDL (Figure 6), we found that they correlate for words, that is, DDL and LDL generally agreed on which subjects show an effect of trial-to-trial learning as modelled by the dynamic simulation. No such correlation was present for nonwords.
@@
Figure 6: Comparison of AIC differences between dynamic and static simulations for DDL and LDL. Each dot represents one participant. The x-axis denotes the difference in AIC between dynamic and static simulations for the LDL models and the y-axis for the DDL models. Negative difference values indicate that the dynamic simulation outperforms the static simulation in terms of AIC, while positive difference values indicate the opposite. If a dot falls within a green shaded area, LDL and DDL agree on whether dynamic simulations provide a better model fit than static simulations for the respective participant. For words, the AIC differences are correlated ($r=.48,p=0.001$); for nonwords no correlation is present ($p=0.4715$). Figure adapted from Heitmeier et al., 2025b .
We conclude from this analysis that DDL is not able to model fine-grained trial-to-trial effects as well as LDL. For words, DDL shows an advantage of learning for a similar number of subjects as LDL but the difference between dynamic and static is smaller for DDL than for LDL. At least for words, DDL and LDL largely agree on which subjects show an effect of trial-to-trial learning. A possible reason for the reduced effect of trial-to-trial learning in DDL is that a more accurate initial mapping may lead to less pronounced error feedback in the update step, thus each learning step having a smaller impact on subsequent trials. Moreover, the larger number of weights in the DDL networks might also mean that weight updates are distributed more evenly across weights and not centered as strongly on the weights of cues shared with subsequent words.
@@ 5 Discussion and Conclusion
In this paper we explored whether replacing the linear mappings (Linear Discriminative Learning, LDL) in the Discriminative Lexicon Model (DLM) with dense neural networks (termed Deep Discriminative Learning, DDL) improves the model’s mapping accuracy. More crucially for cognitive modelling however, we also investigated whether such improved mapping accuracy also increases the model’s performance at predicting behavioural data. We found that across four languages (English, Dutch, Taiwan Mandarin, Estonian) DDL always improved training accuracy and also accuracy on held-out data for English on comprehension, the dataset with the largest number of different lemmas. For production, DDL performed better than LDL across both training and test data for all four datasets. When predicting average lexical decision reaction times, we found that a frequency-informed variant of DDL (FIDDL) where the model is trained on each word according to its frequency outperformed all other methods (including frequency-informed LDL, FIL). However, when predicting fine-grained trial-to-trial lexical decision reaction times, DDL performs worse than LDL, even though the two generally agree as to which subjects show a better model fit when based on simulations where the DLM is dynamically updated after each trial.
@@
It is worthwhile to focus on a few aspects of these results in more detail. First, we find a notable difference across datasets: In comprehension, DDL only reliably improves mapping accuracy over that of LDL for English, but not for Mandarin and Estonian. This may to some extent be explainable by the respective data sizes: the English dataset is larger (28,000 word forms) than the Estonian and Mandarin ones (6,000 and 3,000). However, the nature and diversity of the datasets may also play role. The English and Dutch datasets have a much larger number of different lemmas with few inflected word forms per lemma, while the Estonian dataset only has 230 lemmas with a large number of (relatively regular) inflected forms. For Mandarin, tone information contributes to a much larger number of cues (9,644, compare this to the much larger English dataset, which only has 7,030 cues), and as we saw in Section 3, fewer shared cues increase the performance of LDL models in comprehension. This suggests that the more semantically diverse a dataset, the more it profits from modelling with DDL. In production, on the other hand, all four models profit from DDL. In Heitmeier et al., 2025b we found that this can be accounted for to a significant extent by the loss function: we used the binary cross-entropy loss to train a linear model with the same number of trainable parameters as LDL but with an added sigmoid function and found that it achieved almost the same accuracy as the deep learning models. However, the advantage of DDL over LDL in the present setup may also be due to the representations used. Chuang et al., (2024) found that LDL performs surprisingly well at mapping contextualised embeddings to pitch contours, and preliminary results suggest that in this setup, DDL does not improve over LDL. [^7] However, mapping accuracy from form (low-dimensional) to embeddings (high-dimensional) did improve compared to linear mappings, which is perhaps unsurprising given the difference in dimensionality. This indicates that when both input and output are sufficiently discriminable, LDL suffices.
@@
Second, notice that while DDL models outperformed LDL models in terms of pure mapping accuracy at least for the English dataset, they do not necessarily outperform LDL at predicting reaction times: DDL outperforms LDL at predicting average reaction times in the Dutch Lexicon Project (DLP), but not in the British Lexicon Project (BLP). The same is true for modelling lexical decision reaction times at the individual subject-level: even leaving aside trial-to-trial learning, the DDL models never perform better than the LDL models for static simulations. This highlights that pure mapping accuracy is not enough for predicting reaction times. However, it also leads to a further point, namely that modern machine learning and artificial intelligence may in some aspects be too good: human speakers make a lot of mistakes and are not particularly accurate. For example, participants in the DLP agree with themselves at chance level for low frequency words (below 10 per million) when tested on the same words twice (Diependaele et al.,, 2012). In the study of Ramscar et al., (2013), young adults (mean age 21.1 years) performed at chance level for low frequency words in an English lexical decision task (Balota et al.,, 1999). Moreover, the question arises of whether fully optimized performance in the DLM would be in any way helpful for predicting reaction times. Dense neural networks can in principle learn any continuous function to an arbitrary point of accuracy when given enough parameters (Hornik et al.,, 1989). In the case of the DLM this would mean that all predicted semantic and form vectors would be entirely accurate, thus losing any information we can gain from the DLM above and beyond simply using binary form vectors and pre-trained embeddings. The DLM’s imprecisions and mistakes offer valuable insights into possible processing bottlenecks.
@@
Third, examining the words that DDL maps more accurately than LDL is illuminating for linguistic analysis. We found that LDL outperforms DDL for words which contain very unique trigrams (e.g. #px, #oz), while DDL outperforms LDL when words contain trigrams which express a range of different functions such as -er. The re-use of building blocks is ubiquitous in language. While humans are able to deal with this “bricolage” relatively effortlessly, it is nevertheless a processing bottleneck which LDL struggles with. In this way, the comparison between LDL and DDL shows which components of single word processing may be particularly challenging. This finding could also have implications for theories of morphological priming (e.g. Taft and Forster,, 1975; Rastle and Davis,, 2008): our results suggest that LDL may place pseudomorphological words (e.g. barriers, choler) in the semantic neighbourhood of morphologically complex words (e.g. livelier, chunkier). This could explain why morphologically transparent and morphologically opaque words are primed in a similar way (e.g. Creemers et al.,, 2020) even without morphological decomposition (for linear DLM models capturing morphological priming effects, see Baayen and Smolka,, 2020; Chuang et al.,, 2023). Again, however, this result may, at least in part, be due to the representations used. An examination of the final segment of words ending in -er in the Buckeye corpus (Pitt et al.,, 2007) reveals that they have more varied pronunciations in nouns than in adjectives. These results dovetail well with the hypothesis that the nouns, which have much more varied meanings, have more differentiated realizations of the exponent — a distinction that is neutralized in the orthography. This in turn suggests that deep learning is required for -er words in English because the representations are impoverished, which opens the possibility that a linear mapping based on actual articulation might suffice.
@@
Fourth, our results highlight the importance of the data distribution the models are trained on for cognitive modelling. Using a training/validation/test split is not a plausible mechanism for training cognitive models: after all, we usually encounter high frequency (seen) words, with the occasional low frequency (possibly unseen) word interspersed in-between. FIDDL (frequency-informed DDL) therefore has a much more plausible learning setup: it is trained on the full token distribution. By calculating the model’s accuracy across all words in the dataset we get a more realistic picture of how the model would fare with a real-world distribution of tokens (see also Heitmeier et al., 2023a, , for a discussion of token-based accuracy calculations). Surprisingly, this accuracy is very high (97% and 87% for Dutch and English respectively) even though the model saw 40% of the words in the English and 9% of the words in the Dutch dataset only 10 times or fewer. Such a setup also quells concerns about the performance of computational models on held-out data (Pinker and Ullman,, 2002). Humans are able to generalise to unseen data, but in reality, word distributions are skewed towards high frequency words. Evidently, both humans and FIDDL perform excellently in such a setting. Nonetheless, one might argue that particularly in languages with rich morphology such as Finnish, speakers may still need to generalise to low frequency unseen word forms regularly. As shown by Nikolaev et al., (2024) for a set of some 50,000 Finnish inflected nouns, EL can already perform very well on both training data and held-out low-frequency words. To address this question with further precision, a natural next step is to evaluate performance of FIDDL on these and similar data sets. In any case, these results once again highlight the importance of training models of lexical processing in a frequency-informed way when one is interested in a usage-based analysis of the data. The frequency-informed FIDDL and FIL models by far outperform their non-frequency informed counterparts when predicting average reaction times in the BLP and DLP.